OpenStudy (death47):

What is What is -12 / (2/3)?\ A.-36 B.-24 C.-18 D.-8 I think its C.

OpenStudy (death47):

@Awolflover1 @savana24

OpenStudy (marihelenh):

\[\frac{ -12 }{ \frac{ 2 }{ 3 } }\] is the same as \[-12 * \frac{ 3 }{ 2 }\]

OpenStudy (death47):

yea

OpenStudy (marihelenh):

So, you were correct, it is C.